# Enhancing Retrosynthesis with Conformer: A Template-Free Method

### v1

@@ Abstract
Retrosynthesis plays a crucial role in the fields of organic synthesis and drug development, where the goal is to identify suitable reactants that can yield a target product molecule. Although existing methods have achieved notable success, they typically overlook the 3D conformational details and internal spatial organization of molecules. This oversight makes it challenging to predict reactants that conform to genuine chemical principles, particularly when dealing with complex molecular structures, such as polycyclic and heteroaromatic compounds. In response to this challenge, we introduce a novel transformer-based, template-free approach that incorporates 3D conformer data and spatial information. Our approach includes an Atom-align Fusion module that integrates 3D positional data at the input stage, ensuring correct alignment between atom tokens and their respective 3D coordinates. Additionally, we propose a Distance-weighted Attention mechanism that refines the self-attention process, constricting the model’s focus to relevant atom pairs in 3D space. Extensive experiments on the USPTO-50K dataset demonstrate that our model outperforms previous template-free methods, setting a new benchmark for the field. A case study further highlights our method’s ability to predict reasonable and accurate reactants.
@@ 1 Introduction
Retrosynthesis, initially formulated by Corey (1991), stands as a fundamental problem within organic synthesis, aiming to discover and predict reactants given product molecules. This task presents a large chemical search space and myriad molecular combinations, posing significant challenges. Traditionally, researchers relied heavily on extensive knowledge of organic chemistry and reaction mechanisms, often resulting in inefficiency and heavy dependency on expertise. In recent years, with the significant impact of data-driven deep learning methods across various fields, computer-aided Retrosynthesis has gained increasing attention, especially the template-free methods.
@@
Existing methods are categorized into three types based on reliance on template prior: template-based, semi-template, and template-free. Template-based methods Chen and Jung (2021); Coley et al. (2017); Dai et al. (2019) view Retrosynthesis as a template retrieval task and subsequently employ the retrieved templates (specific chemical changes) to transform the input product molecule into reactants using cheminformatics tools like RDKit Landrum and others (2013). Semi-template methods Shi et al. (2020); Yan et al. (2020) follow two steps: i) Reaction Center Identification: identifying the reaction centers of the product using product-reactant aligned atom mapping and then using RDKit to convert the product into synthons; ii) Synthons Completion: building another model to complete synthons into reactants. Template-free methods regard molecules as sequences, like the Simplified Molecular Input Line Entry System (SMILES) Weininger (1988), and frame Retrosynthesis prediction as machine translation. So that it can implicitly learn reaction rules and easily scale to larger data sets without relying on external template databases or molecular editing with RDKit.
@@
Problem. In previous studies, the typical representation of molecules involves 1D sequences (i.e., SMILES) and 2D graphs. They overlooked the perception and comprehension of 3D conformer information, which is crucial for understanding molecular stereochemistry, reaction centers, chirality, and other properties Meng et al. (2023); Dang et al. (2018). Especially in molecules with intricate 3D structures (e.g., polychiral, heteroaromatic), the absence of 3D conformer information can lead to synthesis predictions that defy established chemical rules.
@@
In this work, we focus on template-free Transformer for Retrosynthesis prediction. Incorporating 3D conformer information into this framework presents two challenges: i) integrating 3D features while maintaining alignment between atom tokens and corresponding 3D representations; ii) controlling receptive field of model based on spatial structure. To tackle these two challenges, we propose our method, a novel end-to-end Retrosynthesis Transformer that encodes both 1D SMILES sequences and 3D conformer information. We propose two modules, Atom-align Fusion and Distance-weighted Attention, to overcome the aforementioned challenges respectively. The model can jointly encode the 1D sequential and 3D positional information of molecules and redistribute attention weights based on 3D distances between atoms. Moreover, data augmentation Tetko et al. (2020); Zhong et al. (2022), like shuffling the atom order or aligning root atom between product and reactants, are employed to enhance the overall performance. Finally, we introduce SMILES alignment Deshpande and Narasimhan (2020); Wan et al. (2022) by incorporating attention guidance loss during training. Our model introduces 3D conformer information, allowing it to generate reasonable synthesis predictions even for molecules with intricate 3D structures, and it does not rely on additional molecular editing. We conducted experiments using a widely recognized dataset USPTO-50k Schneider et al. (2016). Results demonstrate that our model can enhance performance in terms of Top-k accuracy in both known and unknown reaction classes. Our proposed method achieves a new state-of-the-art for template-free methods, and it is also comparable to template-based and semi-template-based methods. Our contributions are summarized as follows:
  • To the best of our knowledge, the proposed model our method is the first to integrate 3D conformer information into template-free method and train the model in an end-to-end manner.
@@
  • Experiments on the USPTO-50k dataset indicate that proposed model our method achieves state-of-the-art Retrosynthesis performance in template-free methods with a Top-1 accuracy of 55.5% and 67.2% without and with reaction class, respectively.
@@ 2 Related Work
Template-based methods utilize reaction templates from pre-constructed database to capture the rules of chemical changes and match input products with the templates. RetroSim Coley et al. (2017) is a similarity-based approach using molecular fingerprint similarity to rank candidate templates. GLN Dai et al. (2019) employs the Conditional Graph Logic Network to learn when rules from reaction templates should be applied, implicitly considering chemical feasibility and strategy. Inspired by the chemical intuition that molecular changes primarily occur locally during reactions, LocalRetro Chen and Jung (2021) is locally encoded and refined to accommodate non-local reactions. Although these methods perform the best among the three categories, their reliance on pre-defined templates limits their coverage and scalability, rendering them insufficient for real-world scenarios.
Semi-template methods do not directly use templates but instead combine chemical rules with generative models. Existing methods typically follow a two-stage procedure: i) reaction center identification; ii) synthons completion. Reaction center identification predicts the reaction center of the product and breaks the corresponding bond into incomplete molecules, called synthons. Synthons completion uses generative models to complete synthons into reactants. RetroXpert Yan et al. (2020) builds an Edge-enhanced Graph Attention Network (EGAT) that takes a molecule graph as input and predicts bond disconnection probabilities to obtain synthons. It then employs a seq2seq model to generate reactants in SMILES format. In contrast to the SMILES format, G2Gs Shi et al. (2020) completes synthons into reactants in graph format through node selection and edge labeling. GraphRetro Somnath et al. (2020) expands synthons into complete reactants by attaching relevant leaving groups. RetroPrime Wang et al. (2021) introduces a two-stage procedure by adding additional labels to SMILES. In contrast, MEGAN Nijssen and Kok (2004) and Graph2Edits Zhong et al. (2023) reframe the generative process as a series of graph edits executed by RDKit.
@@
Figure 1: Architecture overview of our method based on Transformer. a) SMILES and molecular conformer are processed through Word Embedding and ComENet to yield SMILES sequence embedding and 3D position embedding. Following the Atom-align integration, fusion embedding is then input into the Encoder. b) In Encoder, 3D Distance weight is obtained for redistribution of self-attention by Gaussian Basis Kernel function, while also refined layer-by-layer. c) In Decoder, SMILES Alignment is employed as guidance to align the cross-attention between the product and reactants.
@@ 3.1 Template-free Retrosynthesis
Molecules can be represented as sequences of tokens in SMILES format. Following the SMILES tokenization of Schwaller et al. (2019) that separates each atom (i.e., O, C, N, Cl) and non-atom tokens (i.e., -, =, #), we let $S=[s_{1},s_{2},\dots s_{n}]$ denote SMILES sequence with $n$ number of tokens. In template-free Retrosynthesis, we obtain the product SMILES $S_{P}$ and the reactants SMILES $S_{R}$, and attempt to find a method $f_{\theta}$ that can map $S_{P}$ to $S_{R}$ with auto-regressive way: $S_{R}=f_{\theta}(s_{i+1}|s_{0},\dots,s_{i},S_{P})$ in $i$ step.
@@ 3.2 Molecular Conformer and ComENet
Molecule is a dynamic structure since atoms are in continual motion in 3D space. The local minima on the potential energy surface are called conformers. Generally, molecular conformer can be represented as $G_{\text{3D}}=(A,C)$, where $A=\{a_{i}\}_{i=1,\dots,n}$ is atoms list, $C=\{(x_{i},y_{i},z_{i})\}_{i=1,\dots,n}$ is a 3D coordinate list. To fulfill the global completeness of 3D conformer information of a molecule, ComENet Wang et al. (2022) encodes rotation angles and spherical coordinates within the 1-hop neighborhood of atoms into the message passing scheme Balcilar et al. (2021).
@@ 4 Method
In this section, we introduce our method, a novel transformer-based model that is capable of learning richer molecular representation through SMILES sequence and spatial conformer. An overview of the proposed method is shown in Figure 1.
Figure 1a illustrates Atom-align Fusion that incorporates both 1D SMILES tokens embedding and 3D position embedding to get fusion embedding as inputs for encoder-decoder model. By adopting this approach, the model is capable of integrating 3D positional information while preserving alignment (Section 4.1). Figure 1b illustrates Distance-weighted Attention, which directs self-attention to prioritize chemical relevant atoms within spatial structure (Section 4.2). Figure 1c illustrates SMILES alignment strategy for further bridging the mapping relationship between product SMILES and reactant SMILES in decoder cross-attention (Section 4.3). The overall training and inference process is conducted in an end-to-end manner, and it is a template-free method that does not rely on RDKit for molecular editing.
@@ 4.1 Atom-align Fusion
The most direct method of feature integration is concatenation, but this approach disrupts the alignment between 1D representations of atom tokens and their corresponding 3D representations. The challenge lies in how to integrate conformer information into the molecular representation while preserving alignment between them.
@@
Firstly, to obtain the complete 3D molecular representation, we extracted 3D position embedding $P_{\text{3D}}$ from the molecular conformer using ComENet Wang et al. (2022). ComENet proposes a novel message passing scheme Balcilar et al. (2021) that operates within a 1-hop neighborhood with radial distance $d$, polar angle $\theta$, azimuthal angle $\phi$ and rotation angle $\tau$, and then builds message passing scheme as
$\mathbf{v}_{i}^{t+1}=g\left(\mathbf{v}_{i}^{t},\sum_{j\in\mathcal{N}_{i}}f\left(\mathbf{v}_{j},d_{ij},\theta_{ij},\phi_{ij},\tau_{ij}\right)\right)$ (1)
@@
Subsequently, we input $P_{\text{3D}}$ along with the SMILES tokens embedding $T$ into the encoder. To facilitate a more comprehensive integration of SMILES and conformer, we aligned tokens embedding $T$ and $P_{\text{3D}}$ with atom index in the same vector space. We initially enlarge $P_{\text{3D}}$ from $N$ (number of atoms) length to $M$ (number of tokens) length by padding 0 to the indexes of non-atom tokens. In model input stage, we blend the $P_{\text{3D}}$ and $T$ to get the fusion embedding $F_{\text{3D}}$, allowing effective integration of 3D positional and sequential information at each atom position while preserving the alignment between them:
$F_{\text{3D}}=\lambda_{1}\times\operatorname{pad}(P_{\text{3D}})+\lambda_{2}\times T$ (2)
@@ 4.2 Distance-weighted Attention
Similar to the method of partitioning attention heads in Wan et al. (2022), we divide the multi-head attention into normal attention heads and spatial attention heads. Normal attention heads follow the traditional transformer manner to learn sequence-wide information, whereas spatial attention heads employ 3D distances as an external attention weight to enable attention redistributed based on 3D distances. The whole module can be formulated as two steps: 1) Construct 3D distance weight, 2) Attention Redistribute & Weight Refine.
Construct 3D Distance Weight
First, we construct a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ from molecular conformer of molecule as
$\mathbf{D}=\{\mathbf{d}_{i,j}\}_{i,j=1,2,\dots,N}\quad\mathbf{d}_{ij}=\|\mathbf{c}_{i}-\mathbf{c}_{j}\|_{2}$ (3)
where $N$ is number of atoms, $i,j$ are index of atoms pair, $\mathbf{c}_{i}$ is coordiante of $i$-th atom. To explore high-dimensional space, we employ Gaussian Basis Function to transform the distance $\mathbf{d}_{ij}$ between each atom pair ($i$, $j$) into a higher dimension as
$\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}\left|\sigma^{k}\right|}\exp\left(-\frac{1}{2}\left(\frac{\gamma_{(i,j)}\mathbf{d}_{ij}+\beta_{(i,j)}-\mu^{k}}{\left|\sigma^{k}\right|}\right)^{2}\right)$
where k={1, …, K}, K is the number of Gaussian Kernel kernels. $\mu^{k}$ and $\sigma^{k}$ represent kernel center and scaling factor of $k$-th Gaussian Basis kernel. $\gamma_{(i,j)}$ and $\beta_{(i,j)}$ are learnable scalars related to bond type (i.e., SINGLE, DOUBLE) of pair ($i$, $j$). To obtain the 3D Distance weight $\Phi_{(i,j)}$, we use two layers of Non-Linear Layer with the activation function Gaussian Error Linear Units ($\operatorname{GELU}$) Hendrycks and Gimpel (2016) to encode $\psi_{(i,j)}$ as
$\Phi_{(i,j)}=\mathbf{W_{2}}(\operatorname{GELU}(\mathbf{W_{1}}(\psi_{(i,j)})))$ (5)
@@
Attention Redistribute & Weight Refine
Second, we leverage the previously obtained 3D Distance weight $\psi_{(i,j)}$ to guide the model in redistributing attention by the internal spatial structure, instead of directly calculating attention scores between each token like vanilla Transformer. We divide the attention heads into two halves: normal attention head and spatial attention head. The normal attention head still captures the context of the SMILES sequence like original attention computation. The spatial attention head, on the other hand, focuses more on the molecular internal 3D structure, limiting the receptive field of each atom and considering the 3D distance with other atoms. For the $i$-th token of the $l$-th encoder layer, the roll-out form of spatial attention head is formulated as:
$x_{i,\text{spatial}}^{l+1}$ $=\sum_{j\in N(i)}\operatorname{softmax}\left(\Phi_{(i,j)}^{l}\odot\frac{q_{i}k_{j}^{T}}{\sqrt{d}}\right)v_{j}$ (6)
${\left[q_{i},k_{j},v_{j}\right]}$ $=\left[h_{i}^{l}\mathbf{W^{Q}},h_{j}^{l}\mathbf{W^{K}},h_{j}^{l}\mathbf{W^{V}}\right]$
where $\Phi_{(i,j)}^{l}$ is almost the same as 3D Distance weight $\Phi_{(i,j)}$ but padding 0 in non-atom token for element-wise multiplication $\odot$, $\mathbf{W^{Q}}$, $\mathbf{W^{K}}$ and $\mathbf{W^{V}}$ are projection matrices for query $q$, key $k$, value $v$ in self-attention. Computed representations of normal attention head and spatial attention head are concatenated along the hidden dimension and then processed through a linear layer, forming the updated context for the next layer $h^{l+1}$ as:
$h^{l+1}=\mathbf{W_{Linear}}\left(\left[x^{l+1}_{\text{normal}};x^{l+1}_{\text{spatial}}\right]\right)$ (7)
In addition, we perform Weight Refine to update our 3D Distance weight layer by layer. The Weight Refine module is a fully connected (FC) layer that takes concatenation of updated contexts of atoms pair ($i$, $j$) to refine the 3D attention weight among receptive field:
$\Phi_{(i,j)}^{l+1}=\Phi_{(i,j)}^{l}+\operatorname{FC}\left(\left[h_{i}^{l+1};h_{j}^{l+1}\right]\right)$ (8)
4.3 Training Strategy
SMILES Alignment
During the occurrence of chemical reactions, the majority of atoms remain unchanged. We can utilize the atom-map to annotate the correspondence between atoms of product and reactants in reaction, which remains unchanged during the reaction. Furthermore, the correspondence of atoms can be readily transferred to the token correspondence in the SMILES sequence (see in Figure 1c) and construct SMILES Alignment Map (SAM) as:
[…]

### v2

# Retro3D: A 3D-aware Template-free Method for Enhancing Retrosynthesis via Molecular Conformer Information

@@ Abstract
Retrosynthesis plays a crucial role in organic synthesis and drug discovery, focusing on identifying a set of reactants capable of synthesizing a target product molecule. Although the existing approaches have shown promising results, they often overlook 3D conformer information and molecular spatial structure, limiting their ability to generate reactants that comply with chemical rules, especially for complex molecules (e.g., polychiral, heteroaromatic). To tackle this problem, we propose Retro3D, a transformer-based template-free method that integrates molecular conformer information and spatial structure. We devise an Atom-align Fusion module to combine 3D positional information in model input stage, ensuring alignment between atom tokens and corresponding 3D representations. Furthermore, we design a Distance-weighted Attention mechanism to guide self-attention, constraining receptive field of model and emphasizing chemical relevant atom pairs in 3D space. Experiments conducted on the USPTO-50K and USPTO-FULL datasets demonstrate that Retro3D significantly outperforms existing template-free approaches, achieving a new state-of-the-art performance. Case studies further highlight its capability to predict accurate and chemically plausible reactants, even for target molecules with intricate structures.
@@ 1 Introduction
Retrosynthesis, initially formulated by Corey (1991), stands as a fundamental problem within organic synthesis, aiming to discover and predict reactants given product molecules. This task presents a large chemical search space and myriad molecular combinations, posing significant challenges. Traditionally, researchers relied heavily on extensive knowledge of organic chemistry and reaction mechanisms, often resulting in inefficiency and heavy dependency on expertise. In recent years, with the significant impact of data-driven deep learning methods across various fields, computer-aided Retrosynthesis has gained increasing attention, especially the template-free methods Zhong et al. (2024).
@@
Existing methods are categorized into three types based on reliance on template prior: template-based, semi-template, and template-free. Template-based methods Chen and Jung (2021); Coley et al. (2017); Dai et al. (2019) view Retrosynthesis as a template retrieval task and subsequently employ the retrieved templates (specific chemical changes) to transform the input product molecule into reactants using cheminformatics tools like RDKit Landrum and others (2013). Semi-template methods Shi et al. (2020); Yan et al. (2020) follow two steps: i) Reaction Center Identification: identifying the reaction centers of the product using product-reactant aligned atom mapping and then using RDKit to convert the product into synthons; ii) Synthons Completion: building another model to complete synthons into reactants. Template-free methods Zheng et al. (2019); Tetko et al. (2020); Mao et al. (2021) regard molecules as sequences, like the Simplified Molecular Input Line Entry System (SMILES) Weininger (1988), and frame Retrosynthesis prediction as machine translation. So that it can implicitly learn reaction rules and easily scale to larger data sets without relying on external template databases or molecular editing with RDKit.
@@
Problem. In previous studies, the typical representation of molecules involves 1D sequences (i.e., SMILES) and 2D graphs. They overlooked the perception and comprehension of 3D conformer information, which is crucial for understanding molecular stereochemistry, reaction centers, chirality, and other properties Meng et al. (2023); Dang et al. (2018). Especially in molecules with intricate 3D structures (e.g., polychiral, heteroaromatic), the absence of 3D conformer information can lead to synthesis predictions that defy established chemical rules Imrie et al. (2020).
@@
In this work, we focus on template-free Transformer for Retrosynthesis prediction. Incorporating 3D conformer information into this framework presents two challenges: i) integrating 3D features while maintaining alignment between atom tokens and corresponding 3D representations; ii) constraining receptive field of model based on spatial structure. To tackle these two challenges, we propose Retro3D, a novel end-to-end Retrosynthesis Transformer that encodes both 1D SMILES sequences and 3D conformer information. We propose two modules, Atom-align Fusion and Distance-weighted Attention, to overcome the aforementioned challenges respectively. The model can jointly encode the 1D sequential and 3D positional information of molecules and redistribute attention weights based on 3D distances between atoms. Moreover, data augmentation Tetko et al. (2020); Zhong et al. (2022), like shuffling the atom order or aligning root atom between product and reactants, are employed to enhance the overall performance. Finally, we introduce SMILES alignment Deshpande and Narasimhan (2020); Wan et al. (2022) by incorporating attention guidance loss during training. Our model introduces 3D conformer information, allowing it to generate reasonable synthesis predictions even for molecules with intricate 3D structures, and it does not rely on additional molecular editing. We conducted experiments using two widely recognized datasets: USPTO-50k Schneider et al. (2016), USPTO-FULL. Results demonstrate that our model can enhance performance in terms of Top-k accuracy in both known and unknown reaction classes. Our proposed method achieves a new state-of-the-art for template-free methods, and it is also comparable to template-based and semi-template-based methods. Our contributions are summarized as follows:
  • We propose Retro3D, an end-to-end trained, template-free method that integrates 3D conformer information to generate synthesis routes more in line with real-world scenarios for intricate molecular structures.
@@
  • Experimental results on the USPTO-50k and USPTO-FULL datasets demonstrate that the proposed model, Retro3D, achieves state-of-the-art retrosynthesis performance among template-free methods, under both with and without reaction class conditions.
@@ 2 Related Work
Template-based methods utilize reaction templates from pre-constructed database to capture the rules of chemical changes and match input products with the templates. RetroSim Coley et al. (2017) is a similarity-based approach using molecular fingerprint similarity to rank candidate templates. GLN Dai et al. (2019) employs the Conditional Graph Logic Network to learn when rules from reaction templates should be applied, implicitly considering chemical feasibility and strategy. Inspired by the chemical intuition that molecular changes primarily occur locally during reactions, LocalRetro Chen and Jung (2021) is locally encoded and refined to accommodate non-local reactions. Although these methods have strong interpretability, their reliance on pre-defined templates limits their coverage and scalability, rendering them insufficient for real-world scenarios.
Semi-template methods do not directly use templates but instead combine chemical rules with generative models. Existing methods typically follow a two-stage procedure: i) reaction center identification; ii) synthons completion. Reaction center identification predicts the reaction center of the product and breaks the corresponding bond into incomplete molecules, called synthons. Synthons completion uses generative models to complete synthons into reactants. RetroXpert Yan et al. (2020) builds an Edge-enhanced Graph Attention Network (EGAT) that takes a molecule graph as input and predicts bond disconnection probabilities to obtain synthons. It then employs a seq2seq model to generate reactants SMILES. In contrast to SMILES format, G2Gs Shi et al. (2020) completes synthons into reactants in graph format through node selection and edge labeling. GraphRetro Somnath et al. (2020) expands synthons into complete reactants by attaching leaving groups. RetroPrime Wang et al. (2021) introduces a two-stage procedure by adding additional labels to SMILES. In contrast, MEGAN Sacha et al. (2021) and Graph2Edits Zhong et al. (2023) reframe the generative process as a series of graph edits executed by RDKit.
@@ 3.1 Template-free Retrosynthesis
Molecules can be represented as sequences of tokens in SMILES format. Following the SMILES tokenization of Schwaller et al. (2019) that separates each atom (i.e., O, C, N, Cl) and non-atom tokens (i.e., -, =, #), we let $S=[s_{1},s_{2},\dots s_{n}]$ denote SMILES sequence with $n$ tokens. In template-free Retrosynthesis, we obtain the product SMILES $S_{P}$ and the reactants SMILES $S_{R}$, and attempt to find a method $f_{\theta}$ that can map $S_{P}$ to $S_{R}$ in auto-regressive way: $S_{R}=f_{\theta}(s_{i+1}|s_{0},\dots,s_{i},S_{P})$ in $i$ step.
@@ 3.2 Molecular Conformer and ComENet
Molecule is a dynamic structure since atoms are in continual motion in 3D space. The local minima on the potential energy surface are called conformers. Generally, molecular conformer can be represented as $G_{\text{3D}}=(A,C)$, where $A=\{a_{i}\}_{i=1,\dots,n}$ is atoms list, $C=\{(x_{i},y_{i},z_{i})\}_{i=1,\dots,n}$ is 3D coordinates list. To fulfill the global completeness of 3D conformer information of a molecule, ComENet Wang et al. (2022) encodes rotation angles and spherical coordinates within the 1-hop neighborhood of atoms via the message passing scheme Balcilar et al. (2021).
Figure 1: Architecture overview of Retro3D based on Transformer. a) SMILES and molecular conformer are processed through Word Embedding and ComENet to yield SMILES sequence embedding and 3D position embedding. Following the Atom-align integration, fusion embedding is then input into the Encoder. b) In Encoder, 3D Distance weight is obtained for redistribution of self-attention by Gaussian Basis Kernel function, while also refined layer-by-layer. c) In Decoder, SMILES Alignment is employed as guidance to align the cross-attention between the product and reactants.
@@ 4 Method
In this section, we introduce Retro3D, a novel transformer-based model that is capable of learning richer molecular representation through SMILES sequence and spatial conformer. An overview of the proposed method is shown in Figure 1.
Figure 1a illustrates Atom-align Fusion that incorporates both 1D SMILES tokens embedding and 3D position embedding to get fusion embedding as inputs for encoder-decoder model. By adopting this approach, the model is capable of integrating 3D positional information while preserving alignment (Section 4.1). Figure 1b illustrates Distance-weighted Attention, which directs self-attention to prioritize chemical relevant atoms within spatial structure (Section 4.2). Figure 1c illustrates SMILES alignment strategy for further bridging the mapping relationship between product SMILES and reactant SMILES in cross-attention module of decoder (Section 4.3). The overall training and inference process is conducted in an end-to-end manner, and it is a template-free method that does not rely on RDKit for molecular editing.
@@ 4.1 Atom-align Fusion
The most direct method of feature integration is concatenation Mao et al. (2021), but this approach disrupts the alignment between 1D representations of atom tokens and their corresponding 3D representations. The challenge lies in how to integrate conformer information into the molecular representation while preserving alignment between them.
@@
Firstly, to obtain the complete 3D molecular representation, we extracted 3D position embedding $P_{\text{3D}}$ from the molecular conformer using ComENet Wang et al. (2022). ComENet proposes a novel message passing scheme Balcilar et al. (2021) that operates within a 1-hop neighborhood with radial distance $d$, polar angle $\theta$, azimuthal angle $\phi$ and rotation angle $\tau$, and then builds message passing scheme as:
$\mathbf{v}_{i}^{t+1}=g\left(\mathbf{v}_{i}^{t},\sum_{j\in\mathcal{N}_{i}}f\left(\mathbf{v}_{j},d_{ij},\theta_{ij},\phi_{ij},\tau_{ij}\right)\right),$ (1)
@@
Subsequently, we input $P_{\text{3D}}$ along with the SMILES tokens embedding $T$ into the encoder. To facilitate a more comprehensive integration of SMILES and conformer, we aligned embeddings $T$ and $P_{\text{3D}}$ with atom index in the same vector space. We initially enlarge the length $P_{\text{3D}}$ from $N$ (number of atoms) to $M$ (number of tokens) by padding 0 to the indexes of non-atom tokens. In model input stage, we blend the $P_{\text{3D}}$ and $T$ to get the fusion embedding $F_{\text{3D}}$, allowing effective integration of 3D positional and sequential information at each atom position while preserving the alignment between them:
$F_{\text{3D}}=\lambda_{1}\times\operatorname{pad}(P_{\text{3D}})+\lambda_{2}\times T,$ (2)
@@ 4.2 Distance-weighted Attention
Similar to the method of partitioning attention heads in Wan et al. (2022), we divide the multi-head attention into normal attention heads and spatial attention heads. Normal attention heads follow the traditional transformer manner to learn sequence-wide information, whereas spatial attention heads employ 3D distances as an external attention weight to enable attention redistributed based on distances. The whole module can be formulated as two steps: 1) Construct 3D distance weight, 2) Attention Redistribute & Weight Refine.
Construct 3D Distance Weight.
First, we construct a distance matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ from molecular conformer of molecule as:
$\mathbf{D}=\{\mathbf{d}_{i,j}\}_{i,j=1,2,\dots,N}\quad\mathbf{d}_{ij}=\|\mathbf{c}_{i}-\mathbf{c}_{j}\|_{2},$ (3)
where $N$ is number of atoms, $i,j$ are indexes of atoms pair, $\mathbf{c}_{i}$ is coordiante of $i$-th atom. To explore high-dimensional space, we employ Gaussian Basis Function Scholkopf et al. (1997) to transform the distance $\mathbf{d}_{ij}$ between each atom pair ($i$, $j$) into a higher dimension as:
$\psi_{(i,j)}^{k}=-\frac{1}{\sqrt{2\pi}\left|\sigma^{k}\right|}\exp\left(-\frac{1}{2}\left(\frac{\gamma_{(i,j)}\mathbf{d}_{ij}+\beta_{(i,j)}-\mu^{k}}{\left|\sigma^{k}\right|}\right)^{2}\right),$
where kk={1, …, K}, K is the number of Gaussian Kernel kernels. $\mu^{k}$ and $\sigma^{k}$ represent kernel center and scaling factor of $k$-th Gaussian Basis kernel. $\gamma_{(i,j)}$ and $\beta_{(i,j)}$ are learnable scalars related to bond type (e.g., SINGLE, DOUBLE) of atom pair ($i$, $j$). To obtain the 3D Distance weight $\Phi_{(i,j)}$, we use two layers of Non-Linear Layer with the activation function Gaussian Error Linear Units ($\operatorname{GELU}$) Hendrycks and Gimpel (2016) to encode $\psi_{(i,j)}$ as:
$\Phi_{(i,j)}=\mathbf{W_{2}}(\operatorname{GELU}(\mathbf{W_{1}}(\psi_{(i,j)}))),$ (5)
@@
Attention Redistribute & Weight Refine.
Second, we leverage the previously obtained 3D Distance weight $\psi_{(i,j)}$ to guide the model in redistributing attention weights by the internal spatial structure, instead of directly calculating attention scores between each token like vanilla Transformer. We divide the attention heads into two halves: normal attention head and spatial attention head. The normal attention head still captures the context of the SMILES sequence like original attention computation. The spatial attention head, on the other hand, focuses more on the molecular internal structure, limiting the receptive field of each atom and considering the 3D distance with other atoms. For the $i$-th token of the $l$-th encoder layer, the roll-out form of spatial attention head is:
$x_{i,\text{spatial}}^{l+1}$ $=\sum_{j\in N(i)}\operatorname{softmax}\left(\Phi_{(i,j)}^{l}\odot\frac{q_{i}k_{j}^{T}}{\sqrt{d}}\right)v_{j},$ (6)
${\left[q_{i},k_{j},v_{j}\right]}$ $=\left[h_{i}^{l}\mathbf{W^{Q}},h_{j}^{l}\mathbf{W^{K}},h_{j}^{l}\mathbf{W^{V}}\right],$
where $\Phi_{(i,j)}^{l}$ is almost the same as 3D Distance weight $\Phi_{(i,j)}$ but padding 0 in non-atom token for element-wise multiplication $\odot$, $\mathbf{W^{Q}}$, $\mathbf{W^{K}}$ and $\mathbf{W^{V}}$ are projection matrices for query $q$, key $k$, value $v$ in self-attention. Computed representations of normal attention head and spatial attention head are concatenated along the hidden dimension and then processed through a linear layer, forming the updated context for the next layer $h^{l+1}$, formally we have:
$h^{l+1}=\mathbf{W_{Linear}}\left(\left[x^{l+1}_{\text{normal}};x^{l+1}_{\text{spatial}}\right]\right).$ (7)
In addition, we perform Weight Refine to update our 3D Distance weight layer by layer. The Weight Refine module is a fully connected (FC) layer that takes concatenation of updated contexts of atoms pair ($i$, $j$) to refine the 3D attention weight among receptive field, which is formulated as:
$\Phi_{(i,j)}^{l+1}=\Phi_{(i,j)}^{l}+\operatorname{FC}\left(\left[h_{i}^{l+1};h_{j}^{l+1}\right]\right).$ (8)
4.3 SMILES Alignment
During the occurrence of chemical reactions, the majority of atoms remain unchanged Chen and Jung (2021); Zhong et al. (2022). We can utilize the atom-map to annotate the correspondence between atoms of product and reactants in reaction, which remains unchanged during the reaction. Furthermore, the correspondence of atoms can be readily transferred to the token correspondence in the SMILES sequence (see in Figure 1c) and construct SMILES Alignment Map (SAM) as:
$SAM_{ij}=\begin{cases}1&\text{ if }R_{i}\underset{\text{ mapped }}{\longleftrightarrow}P_{j}\\ 0&\text{ else }\end{cases}$ (9)